# Planning vs Reasoning: Ablations to Test Capabilities of LoRA layers

### v1

@@ Abstract
Low-Rank Adaptation (LoRA) layers have emerged as a promising approach for efficient model fine-tuning, but their capabilities and limitations have not been fully explored. This paper: 1) Investigates the fundamental question of whether LoRA layers are effective at increasing reasoning + planning abilities 2) We introduce HashChain Reasoning, a novel evaluation dataset that deterministically tests reasoning capabilities.
Through systematic ablation studies on GPT-2, we demonstrate that reasoning capabilities appear to exist primarily in low-rank spaces and can be effectively enhanced using LoRA layers. The effective rank analysis of trained LoRA matrices reveals a 2-3x lower rank requirement for reasoning tasks compared to planning tasks, giving context on where LoRA layers would be effective. This also provides evidence for reasoning fundamentally preferring low-parameter spaces for generalization.
@@ 1 Background
Efficiently handling continual learning for new skills is essential to scale language models [10]. One current approach that has gained traction in the past year is Low Rank Adaptation (LoRA) layers, which add a couple of low rank linear parameters to efficiently fine tune the model [13]. These parameters can either be applied dynamically or merged into the model via a linear combination. An important question here is the learning capacity of LoRA layers and where the boundaries might be.
@@ 1.1 Continual Learning
Replay-based methods use a buffer of past actions which can be used in conjunction with fine tuning to make sure the model doesn’t forget key attributes [10]. With reasoning being an emergent property without any direct training data, this proves to be hard. Regularization is also utilized here but seems to be a half-fix without giving guarantees on knowledge retention and minimizes the effectiveness of finetuning [10].
@@ 1.2 Low-Rank Adaptation (LoRA) Layers
Low-Rank Adaptation (LoRA) represents a significant advancement in parameter-efficient fine-tuning methodologies [6]. By introducing low-rank decomposition matrices into the adaptation process, LoRA enables model specialization while utilizing only a fraction of the parameters typically required for full fine-tuning—replacing an $(n,m)$ matrix with two $(n,rank)\times(rank,m)$ matrices.
This parameter efficiency leads to reduced catastrophic forgetting and minimizes distribution shift during adaptation [3]. The approach is particularly compelling from a theoretical perspective, as the constrained parameter space provides implicit regularization against overfitting [13]. This aligns closely with the minimum description length principle-by restricting the dimensionality of possible adaptations, we can identify minimal distribution shifts that achieve task-specific optimization while preserving the model’s core capabilities.
Other methods include mixture of expert models, and training new experts for tasks [11]. This seems promising but LoRA layers have benefits in portability, training, and inference capacity [12].
@@ 2 Approach
For the experimental setup GPT-2 was utilized alongside a custom training script for the LoRA layer addition [8]. First the model is fine-tuned to its capacity on a certain set of reasoning evals (we will get into what these evals are later). Then using LoRA module, we can try to increase the model’s capabilities on these evals. If possible, this suggests reasoning/circuits related to reasoning are low-rank and could be augmented to increase abilities or for cross domain transfer. Further analysis on the LoRA layers themselves could be done to determine the functional rank of the matrices.
Reasoning and planning are thought of in terms of tree search algorithms here. Planning would be the depth to which the algorithm can search to and reasoning would be the heuristic or communication between branches. Both are needed for more fundamentally complex models.
@@ 2.1 HashHop Dataset
The first dataset used was HashHop, an eval metric initially created for context window evaluations [5] [7]. A markov chain of hashes is generated and the model is asked to predict n hops ahead of the chain. Hashes are utilized because they are completely random.
@@
Figure 3: Using the same metrics, Shannon entropy [9] (left) and cutoffs (right), there is about a 2x-3x decrease in effective rank for the HashChain reasoning task compared to the baseline HashHop task.
@@ 2.3 Training
The goal is to fully finetune a small model on the eval dataset and see if through LoRA layers our metrics increase in a substantial way. The hyper-parameters were fixed and iterations were done till the loss plateaued at an equilibrium. All training scripts are open source https://github.com/neelr/lora-reasoning-experiments.
@@ 3 Experimental Results
Model Base (fit) Post-LoRA
HashHop 0.283 0.302
Reasoning-3chain 0.391 0.452
Reasoning-4chain 0.192 0.369
Figure 4: Calculating the effective rank of the LoRA matrix trained on the regular HashHop task. Shannon Entropy [9](left) gave a mean of 17.87 and a cutoff (right) gave a mean of 158
@@ 3.1 HashHop Evaluations
The chain length was varied between 1-20, with 15-20 being cut off in the graph for being equally close to random chance. For the regular HashHop evaluations the prediction power increased marginally with LoRA layers within the first couple hops 5(a). Jumps in accuracy started larger,  10% for 2 hop, but quickly deteriorated to random chance.
@@
This suggests that although LoRA layers are helpful for fine-tuning specific aspects like style, it is unable to teach fundamentally new planning capabilities to the network. The jumps in accuracy (2, 3, 4, & 5 hop) were on tasks that the model had already expressed prior functionality in. In Figure 5(b) and 4, ablations were done on the rank of the resultant LoRA matrix. The effective rank of each LoRA matrix seems to hover around 150, suggesting that the task is hard to learn and high rank [9].
The result is also supported in the literature. Due to LLM’s being next token predictors, long term planning is especially hard [1]. If thought of in the same way as the breadth-first HashChain, this proves that vertically scaling reasoning is fundamentally high rank/difficult.
@@ 3.2 HashChain Reasoning Evaluations
Spiking near the end for both could also be attributed to latter layers often needing more LoRA parameters [4].
@@ 4 Conclusion
The 4-chain abilities in the HashChain Reasoning dataset saw large increases in accuracy, suggesting not just improving reasoning but also horizontally adding new capabilities. Distinguishing between planning & reasoning is necessary because as we add LoRA layers, its important to check whether or not LoRA layers are the right fit for the job—especially with continual learning.
Planning ability did increase but only in areas that the model already excelled in. This suggests LoRA layers could be used to “squeeze out” more latent planning ability but planning still seems fundamentally hard for LoRA layers [2].
@@ 4.1 Future Research Directions
The use of effective rank + targeted datasets to check the difficulty of tasks for models to learn seems promising, as well as use of LoRA layers to check effective rank of circuits in models. LoRA layers could possibly be used as a replacement for Mixture of Experts or other techniques for increasing parameter counts in a sparse way for reasoning tasks [11].
@@
Code + benchmarks are public and open source @ https://github.com/neelr/lora-reasoning-experiments

### v2

@@ Abstract
Low-Rank Adaptation (LoRA) layers have emerged as a promising approach for efficient model fine-tuning, but their capabilities and limitations remain unexplored. Through systematic ablation studies on GPT-2, we demonstrate that reasoning capabilities exist primarily in low-rank spaces and can be effectively enhanced using LoRA layers, with trained LoRA matrices showing 2-3x lower rank requirements for reasoning tasks compared to planning tasks. We introduce HashChain Reasoning, a novel evaluation dataset for deterministically testing reasoning capabilities, and propose ELoRA (Entropy LoRA), a new adapter architecture that improves reasoning ability while speeding up convergence. Our findings are validated through experiments showing that ELoRA performs approximately 5% better on GSM8k compared to regular LoRAs, along with demonstrating faster convergence rates.
@@ 1 Background
Efficiently handling continual learning for new skills is essential to scale language models (shi_continual_2024, ). One current approach that has gained traction in the past year is Low Rank Adaptation (LoRA) layers, which add a couple of low rank linear parameters to efficiently fine tune the model (wistuba_continual_2023, ). These parameters can either be applied dynamically or merged into the model via a linear combination. An important question here is the learning capacity of LoRA layers and where the boundaries might be.
@@ 1.1 Continual Learning
Replay-based methods use a buffer of past actions which can be used in conjunction with fine tuning to make sure the model doesn’t forget key attributes (shi_continual_2024, ). With reasoning being an emergent property without any direct training data, this proves to be hard. Regularization is also utilized here but seems to be a half-fix without giving guarantees on knowledge retention and minimizes the effectiveness of finetuning (shi_continual_2024, ).
@@ 1.2 Low-Rank Adaptation (LoRA) Layers
Low-Rank Adaptation (LoRA) represents a significant advancement in parameter-efficient fine-tuning methodologies (hu_lora:_2021, ). By introducing low-rank decomposition matrices into the adaptation process, LoRA enables model specialization while utilizing only a fraction of the parameters typically required for full fine-tuning—replacing an $(n,m)$ matrix with two $(n,rank)\times(rank,m)$ matrices.
This parameter efficiency leads to reduced catastrophic forgetting and minimizes distribution shift during adaptation (biderman_lora_2024, ). The approach is particularly compelling from a theoretical perspective, as the constrained parameter space provides implicit regularization against overfitting (wistuba_continual_2023, ). This aligns closely with the minimum description length principle-by restricting the dimensionality of possible adaptations, we can identify minimal distribution shifts that achieve task-specific optimization while preserving the model’s core capabilities.
Other methods include mixture of expert models, and training new experts for tasks (sukhbaatar_branch-train-mix:_2024, ). This seems promising but LoRA layers have benefits in portability, training, and inference capacity (wen_batched_2024, ).
@@ 1.3 Contributions
  • A new ELoRA (Entropy LoRA) adapter that converges faster and performs better on evals using a linear approximation for effective rank.
@@ 2 Approach
For the experimental setup GPT-2 was utilized alongside a custom training script for the LoRA layer addition (radford_language_2019, ). First the model is fine-tuned to its capacity on a certain set of reasoning evals (we will get into what these evals are later). Then using LoRA module, we try to increase the model’s capabilities on these evals. Given improvements, this suggests reasoning/”circuits related to reasoning” are low-rank and could be augmented to increase abilities or for cross domain transfer. Further analysis on the LoRA layers themselves could be done to determine the functional rank of the matrices.
Here, reasoning and planning are thought of in terms of tree search algorithms. Planning would be the depth to which the algorithm can search to and reasoning would be the heuristic or communication between branches. Both are needed for more fundamentally complex models.
@@ 2.1 HashHop Dataset
The first dataset used was HashHop, an eval metric initially created for context window evaluations (hsieh_ruler:_2024, ) (magic_hashhop:_2024, ). A markov chain of hashes is generated and the model is asked to predict n hops ahead of the chain. Hashes are utilized because they are completely random.
@@
Figure 3: Using the same metrics, Shannon entropy (roy_effective_2007, ) (left) and cutoffs (right), there is about a 2x-3x decrease in effective rank for the HashChain reasoning task compared to the baseline HashHop task.
@@ 2.3 Training
The goal is to fully finetune a small model on the eval dataset and see if through LoRA layers our metrics increase in a substantial way. The hyper-parameters were fixed and iterations were done till the loss plateaued at an equilibrium. All training scripts are open source https://github.com/anon/opensource.
@@ 3 Experimental Results
Model Base LoRA ELoRA
HashHop 0.283 0.302 0.303
Reasoning-3chain 0.391 0.452 0.473
Reasoning-4chain 0.192 0.369 0.451
HashHop Regular LoRA Analysis Refer to caption
Figure 4: Calculating the effective rank of the LoRA matrix trained on the regular HashHop task. Shannon Entropy (roy_effective_2007, )(left) gave a mean of 17.87 and a cutoff (right) gave a mean of 158
@@ 3.1 HashHop Evaluations
The chain length was varied between 1-20, with 15-20 being cut off in the graph for being equally close to random chance. For the regular HashHop evaluations the prediction power increased marginally with LoRA layers within the first couple hops 6(a). Jumps in accuracy started larger,  10% for 2 hop, but quickly deteriorated to random chance.
@@
ELoRA Architecture Refer to caption
Figure 5: ELoRA architecture diagram. Similar to LoRA except for a small linear matrix with maximized entropy.
This suggests that although LoRA layers are helpful for fine-tuning specific aspects like style, it is unable to teach fundamentally new planning capabilities to the network. The jumps in accuracy (2, 3, 4, & 5 hop) were on tasks that the model had already expressed prior functionality in. In Figure 6(b) and 4, ablations were done on the rank of the resultant LoRA matrix. The effective rank of each LoRA matrix seems to hover around 150, suggesting that the task is hard to learn and high rank (roy_effective_2007, ).
The result is also supported in the literature. Due to LLM’s being next token predictors, long term planning is especially hard (bachmann_pitfalls_2024, ). If thought of in the same way as the breadth-first HashChain, this proves that vertically scaling reasoning is fundamentally high rank/difficult.
@@ 3.2 HashChain Reasoning Evaluations
Spiking near the end for both could also be attributed to latter layers often needing more LoRA parameters (gao_higher_2024, ).
4 ELoRA Validation Tests
ELoRA Convergence Refer to caption
Figure 7: ELoRA converges faster than LoRA on average when given the same data
There have been many variations to the LoRA architecture, such as taking weight decompositions like DoRA (Weight-Decomposed Low-Rank Adaptation) and QLoRA (Quantized Low Rank Adaptation) both for reasoning and efficiency gains respectively. It has been hard to find a good variation for LoRA’s that converge fast as well as have good reasoning priors.
Here we introduce ELoRA (an entropy LoRA) to validate our analysis in the prior section 5. By having a linear layer prior to the LoRA that maximizes entropy and disentangles the representations, we can utilize the ability of LoRAs to effectively ”pick out” specific representations into a low rank space. This method has been utilized before to prevent representation collapse. (arefin_seq-vcr:_2024, ) (skean_frossl:_2024, )
The ELoRA is structured similarly to a regular LoRA but with an entropy matrix beforehand with a modified loss term. There is no activation function, which allows the adapter to be merged into the weights after training. Learning rates were set to $31e-$ 3 for the matrix and $51e-$ 5 for the regular low-rank matrices and Phi1.5b was utilized for training. (li_textbooks_2023, )
The use of an external matrix could also allow for faster transfer learning convergence through a learned ”high entropy prior”.
For the loss, we utilize a linear entropy equation coined in FroSSL that is equivalent to Rényi $\alpha$-entropy where $\alpha=2$ (skean_frossl:_2024, ). The benefit is that one does not have to calculate the singular values, which is expensive for large matrices. The loss term of the entropy is given below:
$\mathcal{L}_{\text{entropy}}=\log(|Z^{T}Z|_{F}^{2})$
Table 2: GSM8K Validation Accuracy by Adapter
Adapter GSM8K Validation
Phi 1.5 base 33%
LoRA 43%
ELoRA 48%
The results of the experiment were promising and seem to validate the prior sections. ELoRA converged faster than regular LoRAs 7 and consistently scored higher  5-6% higher in GSM8k math evaluations (cobbe_training_2021, ) 2. ELoRA also consistently outperforms LoRAs in HashHop but much more significantly in HashChain where more reasoning is needed 1.
@@ 5 Conclusion
The introduction of ELoRA provides additional evidence for these observations. On the GSM8k benchmark, ELoRA showed a modest but consistent improvement ( 5%) over standard LoRA implementations, with notably faster convergence rates. The entropy component appears to aid in representation learning, by disentangling the representations.
The 4-chain abilities in the HashChain Reasoning dataset saw large increases in accuracy in regular LoRA’s and ELoRA’s, suggesting not just improving reasoning but also horizontally adding new capabilities. Distinguishing between planning & reasoning is necessary because as we add LoRA layers, its important to check whether or not LoRA layers are the right fit for the job—especially with continual learning.
Planning ability did increase but only in areas that the model already excelled in. This suggests LoRA layers could be used to “squeeze out” more latent planning ability but planning still seems fundamentally hard for LoRA layers (bachmann_pitfalls_2024-1, ).
@@ 5.1 Future Research Directions
The use of effective rank + targeted datasets to check the difficulty of tasks for models to learn seems promising, as well as use of LoRA layers to check effective rank of circuits in models. LoRA layers could possibly be used as a replacement for Mixture of Experts or other techniques for increasing parameter counts in a sparse way for reasoning tasks (sukhbaatar_branch-train-mix:_2024, ).
The promising initial results of ELoRA help the reasoning claims in the paper. More comprehensive benchmarking across a broader range of reasoning tasks (such as BIG-Bench, MMLU, and BBH) would help establish the generality of ELoRA’s benefits beyond GSM8k and HashChain Reasoning. Of particular interest is the potential use of the learned entropy matrix as a transferable prior—preliminary results suggest faster convergence.
Testing these benchmarks with new reasoning models such as Deepseek R1 seems interesting, as intermediary layers seem to contain higher entropy than non CoT distilled counterparts (deepseek-ai_deepseek-r1:_2025, ).
@@
Code + benchmarks are public and open source @ https://github.com/anon/opensource